# Dark Energy Survey Year 6 Results: Cell-based Coadds and Metadetection Weak Lensing Shape Catalogue

### v1

@@ Abstract
We present the Metadetection weak lensing galaxy shape catalogue from the six-year Dark Energy Survey (DES Y6) imaging data. This dataset is the final release from DES, spanning 4422 deg2 of the southern sky. We describe how the catalogue was constructed, including the two new major processing steps, cell-based image coaddition and shear measurements with Metadetection. The DES Y6 Metadetection weak lensing shape catalogue consists of 151,922,791 galaxies detected over $riz$ bands, with an effective number density of $n_{\rm eff}$=8.22 galaxies per arcmin2 and shape noise of $\sigma_{\rm e}=0.29$. We carry out a suite of validation tests on the catalogue, including testing for PSF leakage, testing for the impact of PSF modeling errors, and testing the correlation of the shear measurements with galaxy, PSF, and survey properties. In addition to demonstrating that our catalogue is robust for weak lensing science, we use the DES Y6 image simulation suite (Mau, Becker et al., 2025) to estimate the overall multiplicative shear bias of our shear measurement pipeline. We find no detectable multiplicative bias at the roughly half-percent level, with $m=(3.4\pm 6.1)\times 10^{-3}$, at $3\sigma$ uncertainty. This is the first time both cell-based coaddition and Metadetection algorithms are applied to observational data, paving the way to the Stage-IV weak lensing surveys.
@@
gravitational lensing: weak – cosmology: observations
@@ 1 Introduction
Weak gravitational lensing – a small deflection of light due to an intervening gravitational potential between its source and an observer – is a well-established and competitive observational probe to measure how the structure in the Universe has evolved (Amon et al.,, 2022; Secco, Samuroff et al., 2022; Li et al.,, 2023; Dalal et al.,, 2023; Asgari et al.,, 2021). Such analyses measure the tiny statistical patterns in aligned or correlated shapes of a vast number of galaxies. To do so requires high-accuracy measurements of their shapes that have been extensively validated (for a review, see Mandelbaum 2018). As galaxy samples become larger with surveys such as Euclid [^1] http://sci.esa.int/euclid (Laureijs et al. 2011), the Vera C. Rubin Observatory Legacy Survey of Space and Time [^2] http://www.lsst.org (LSST: LSST Science Collaboration et al. 2009; Ivezić et al. 2019), and the Nancy Grace Roman Space Telescope [^3] https://roman.gsfc.nasa.gov (Roman: Spergel et al. 2015), our systematic uncertainty requirements are more stringent, demanding improved shape measurement techniques.
@@
Non-zero $m$ and $c$ terms can be introduced due to imperfect (i) PSF modeling and (ii) shear estimation. During PSF modeling, sources such as inaccurate characterization of detector effects, impurity in stellar samples, and mismodeling of atmospheric effects can cause the misestimation of PSF models, and the PSF size and ellipticity errors result in $m$ and $c$ respectively. Even if PSF models are accurate, misevaluation of the PSF for individual galaxies during shear estimation can lead to significant multiplicative bias. We will explore this further in this paper. Furthermore, in the low signal-to-noise (S/N) regime, the pixel noise causes prominent “noise bias” (e.g., Hirata & Seljak 2003; Kacprzak et al. 2012; Melchior & Viola 2012; Refregier et al. 2012), because shape estimation is a non-linear function of pixel intensities. Other well-known significant shear biases include “model bias” due to the mismatch of assumed and the actual galaxy light profile during shape estimation (e.g., Kacprzak et al. 2014; Melchior et al. 2010; Voigt & Bridle 2010) and “selection bias” that happens when the galaxy selections depend on shear (e.g., Hirata & Seljak 2003; Bernstein & Jarvis 2002).
@@
In order to account for these biases, modern lensing analyses calibrate the shears that have been estimated. In this step, one measures the shear response $\partial e_{i}/\partial\gamma_{j}$, where the measured galaxy ellipticity ($e_{i}$) responds to an applied shear ($\gamma_{j}$). By determining the shear response for an ensemble of galaxies, one can calibrate the aforementioned biases over a wide range of galaxy populations. In the Shear TEsting Programme (STEP; Heymans et al. 2006; Massey et al. 2007) and the GRavitational lEnsing Accuracy Testing challenges (GREAT; Bridle et al. 2010; Kitching et al. 2013; Mandelbaum et al. 2015), the lensing community has dedicated efforts to develop new unbiased shear estimation and calibration algorithms and mitigate/self-calibrate these biases for the Stage-III surveys to satisfy their systematics budgets such as the Hyper-Suprime Cam Subaru Strategic Program [^4] http://hsc.mtk.nao.ac.jp/ssp/ (HSC: Aihara et al. 2018), the Kilo-Degree Survey [^5] http://kids.strw.leidenuniv.nl/ (KiDS: de Jong et al. 2013) and the Dark Energy Survey [^6] http://www.darkenergysurvey.org/ (DES: The Dark Energy Survey Collaboration 2005). We briefly introduce a few more recent algorithms that are relevant to this paper. The first widely utilized technique is Metacalibration (Sheldon & Huff, 2017; Huff & Mandelbaum, 2017). This method artificially applies a shear to each object and numerically measures the response. It can empirically correct for “noise”, “model”, and “selection” biases to a few percent (Sheldon & Huff 2017; Gatti, Sheldon et al. 2021). Additionally, Bayesian Fourier Domain (BFD;  Bernstein & Armstrong 2014; Bernstein et al. 2016a) can independently self-calibrate the biases by constructing templates from deep-field galaxies and building the probability distribution of Fourier-space moments for wide-field galaxies. We also point out that AnaCal, which computes the shear response of basis functions and propagates the shear response of basic modes to derive the shear response of the final observable, has shown the potential to calibrate shear calibration biases to a sub-percent level (Li & Mandelbaum, 2023; Li et al., 2024).
@@
With this progress, a number of challenges remain, which motivate the advancements in this work. In PSF modeling, Meyers & Burchat (2015) identified the differential chromatic refraction (DCR) and the wavelength dependence of seeing as primary drivers of PSF misestimation. These chromatic effects will be especially essential to be accounted for in Stage-IV surveys. Further, they have already been detected in previous DES analyses (Jarvis et al. (2021), Gatti, Sheldon et al. (2021); hereafter GS20). In shear estimation, as the number of single-epoch (SE) images increases for a given region on the sky, naive image coaddition introduces SE image edges into the final coadd image, which produce discontinuities in the coadded PSF. These PSF discontinuities can bias subsequent shape measurements performed on the coadd image (Mandelbaum et al., 2023). This effect motivated the techniques that use galaxies detected in coadd images but with measurements taken from the SE images. Unfortunately, this approach has two main limitations. First, the computational cost of the shear measurement increases linearly with the number of SE images. While the DES has only $\sim$10 SE images overlapping a given point on the sky in a given band, Stage-IV surveys will, in some cases, have $\cal{O}$(100) SE images (LSST Science Collaboration et al., 2009). Second, detecting objects is a weak lensing shear-dependent operation, and ignoring this dependence causes significant biases in shear measurements (e.g., Hirata & Seljak 2003; Kannawadi et al. 2019; Sheldon et al. 2020). Finally, as the depths of surveys increase, the fraction of objects that are blended increases as well. The increased blend fraction coupled with detection biases can cause catastrophic biases in shear measurements for future surveys (Sheldon et al. 2020; hereafter S20).
@@
To address these issues, we have pursued a new approach to shear measurement which proceeds in two stages. First, we employ a cell-based coadding algorithm to reprocess the entire DES survey. A cell-based coadd in this work is defined as a small coadd image, approximately one arcminute on a side, within which none of the input SE images have edges. This kind of coadd has a well-defined PSF model (Mandelbaum et al., 2023; Armstrong et al., 2024) and so is suitable for weak lensing shape measurements. We lose some survey depth, unfortunately, in the process of constructing cell-based coadds since not all input SE images fully overlap a given region on the sky (Armstrong et al., 2024). This trade-off is discussed in more detail below. Second, we run the Metadetection weak lensing shear estimation algorithm over the cell-based coadd images. The application of Metadetection on simulated Rubin data by Sheldon et al. (2023) has demonstrated that Metadetection can produce unbiased shear measurements in the presence of detection and blending at Rubin LSST ten-year depths (LSST Science Collaboration et al., 2009) to a fractional accuracy better than $\sim 0.1\%$. While this level of accuracy is not required for DES Y6 cosmological analyses, we aim for $m$ around 1.3% or smaller. This level, according to the LSST DESC Science Requirements Document (The LSST Dark Energy Science Collaboration et al., 2018), would be sufficient for LSST Year 1 analysis and certainly for DES Year 6 analysis. The previous DES Y3 Metacalibration pipeline (Gatti, Sheldon et al., 2021) produced significant enough biases due to blending to shift the Y3 cosmological contours by $\approx 1\sigma$ (MacCrann, Becker et al., 2022; Amon et al., 2022) if not corrected.
@@
As for additive bias, previous DES analyses have found non-negligible additive bias in DES data (Gatti, Sheldon et al., 2021) that is beyond cosmic variance and above what was measured in image simulations (MacCrann, Becker et al., 2022). The origin of this bias is unknown, and Amon et al. (2022) chose to subtract it from the data before measuring the shear two-point correlation functions. This procedure is valid under the assumption that the bias is actually constant and when the two-point function measurements are restricted to angular scales where the two-point function amplitudes are significantly larger than the additive signal squared. For non-constant additive signals that are driven by PSF modeling and deconvolution, one can use the data themselves to measure the additive bias signals. For the DES Y3 shear catalog, these biases were shown to be small and could thus be safely ignored. Alternatively, one can use the additive bias measurements to directly marginalize over them during the cosmological analysis (Li et al., 2023). For DES Y6, we will employ similar procedures to account for any residual additive biases, and will also attempt to reduce them using new PSF models with chromatic dependence (Schutt et al., 2025).
@@
For the image simulations, we use an updated version of the DES Y3 image simulation code (MacCrann, Becker et al., 2022), to be presented by Mau, Becker et al. 2025. Finally, this paper is further accompanied by Schutt et al. 2025, which describes the upgraded PSF modeling with Piff (PSFs in the Full FOV; Jarvis et al. 2021) that includes color-dependent modeling. This work reports several diagnostic tests on the models created for Y6 data, showing better PSF models in all $griz$ bands compared to Y3 (Jarvis et al., 2021). Finally, the DES collaboration has a second shear catalogue in preparation using the BFD technique (Gatti, Wetzell et al., 2025), and this catalogue will eventually act as a strong cross-check on this work.
@@ 2.1 DES Data
  • The SE images: The $griz$-band calibrated images from the DES processing pipeline, created by DES Data Management (DESDM) (Morganson et al., 2018).
@@ 2.2 Mock Catalogues
In some validation tests of the Metadetection shear catalogue, we require covariance matrices for statistics extracted from complex operations on the data. These covariance matrices need to include both galaxy shape noise and the underlying sample variance contributions from large-scale structure. For these kinds of tests, mock shape catalogues were generated from the shear maps in gravity-only N-body simulations. We make use of CosmogridV1 (Fluri et al., 2022). The dataset is described in Kacprzak et al. (2023). We use 200 independent realisations at the fixed cosmology ($\Omega_{\rm m}=0.26$, $\sigma_{8}=0.84$, $\Omega_{\rm b}=0.0493$, $n_{\rm s}=0.9649$, $h=0.673$). The simulations provide 69 full-sky lens planes between redshift 0 and redshift 3.5 with HEALPix [^8] http://healpix.sourceforge.net (Górski et al., 2005) nside=2048; the lens planes are converted into convergence planes under the Born approximation (e.g., Eq. 2 from Fosalba et al. 2015), and shear planes are obtained from the convergence maps using a full-sky generalisation of the Kaiser & Squires (1993) algorithm (Chang et al.,, 2018; Jeffrey et al.,, 2021). Mock DES Y6 shear fields are obtained by integrating the shear planes assuming preliminary DES Y6 redshift distributions; last, we create mock shear catalogues by sampling the simulated shear fields at the DES Y6 Metadetection galaxy positions and adding DES Y6 shape noise. This last step is performed by randomly rotating the DES Y6 galaxy ellipticities and adding them to the simulated shear field sampled at each galaxy position. Since we can cut four DES Y6 footprints out of each simulation, this procedure provides 800 independent shear catalogues that can be used to infer covariance matrices and perform statistical tests.

### v2

@@ Abstract
We present the Metadetection weak lensing galaxy shape catalogue from the six-year Dark Energy Survey (DES Y6) imaging data. This dataset is the final release from DES, spanning 4422 deg2 of the southern sky. We describe how the catalogue was constructed, including the two new major processing steps, cell-based image coaddition and shear measurements with Metadetection. The DES Y6 Metadetection weak lensing shape catalogue consists of 151,922,791 galaxies detected over rizriz bands, with an effective number density of $n_{\rm eff}$=8.22 galaxies per arcmin2 and shape noise of $\sigma_{\rm e}=0.29$. We carry out a suite of validation tests on the catalogue, including testing for PSF leakage, testing for the impact of PSF modeling errors, and testing the correlation of the shear measurements with galaxy, PSF, and survey properties. In addition to demonstrating that our catalogue is robust for weak lensing science, we use the DES Y6 image simulation suite to estimate the overall multiplicative shear bias of our shear measurement pipeline. We find no detectable multiplicative bias at the roughly half-percent level, with $m=(3.4\pm 6.1)\times 10^{-3}$, at $3\sigma$ uncertainty. This is the first time both cell-based coaddition and Metadetection algorithms are applied to observational data, paving the way to the Stage-IV weak lensing surveys.
@@
gravitational lensing: weak – cosmology: observations, catalogues, techniques: image processing, methods, and techniques
@@ 1 Introduction
Weak gravitational lensing – a small deflection of light due to an intervening gravitational potential between its source and an observer – is a well-established and competitive observational probe to measure how the structure in the Universe has evolved (Amon, Gruen, Troxel, MacCrann, Dodelson, Choi, Doux, Secco, Samuroff, Krause, Cordero, Myles, DeRose, Wechsler, Gatti, Navarro-Alsina, Bernstein, Jain, Blazek, Alarcon, Ferté, Lemos, Raveri, Campos, Prat, Sánchez, Jarvis, Alves, Andrade-Oliveira, Baxter, Bechtol, Becker, Bridle, Camacho, Carnero Rosell, Carrasco Kind, Cawthon, Chang, Chen, Chintalapati, Crocce, Davis, Diehl, Drlica-Wagner, Eckert, Eifler, Elvin-Poole, Everett, Fang, Fosalba, Friedrich, Gaztanaga, Giannini, Gruendl, Harrison, Hartley, Herner, Huang, Huff, Huterer, Kuropatkin, Leget, Liddle, McCullough, Muir, Pandey, Park, Porredon, Refregier, Rollins, Roodman, Rosenfeld, Ross, Rykoff, Sanchez, Sevilla-Noarbe, Sheldon, Shin, Troja, Tutusaus, Tutusaus, Varga, Weaverdyck, Yanny, Yin, Zhang, Zuntz, Aguena, Allam, Annis, Bacon, Bertin, Bhargava, Brooks, Buckley-Geer, Burke, Carretero, Costanzi, da Costa, Pereira, De Vicente, Desai, Dietrich, Doel, Ferrero, Flaugher, Frieman, García-Bellido, Gaztanaga, Gerdes, Giannantonio, Gschwend, Gutierrez, Hinton, Hollowood, Honscheid, Hoyle, James, Kron, Kuehn, Lahav, Lima, Lin, Maia, Marshall, Martini, Melchior, Menanteau, Miquel, Mohr, Morgan, Ogando, Palmese, Paz-Chinchón, Petravick, Pieres, Romer, Sanchez, Scarpine, Schubnell, Serrano, Smith, Soares-Santos, Tarle, Thomas, To, Weller, and DES Collaboration, 2022; Secco, Samuroff, and others, 2022; Li, Zhang, Sugiyama, Dalal, Terasawa, Rau, Mandelbaum, Takada, More, Strauss, Miyatake, Shirasaki, Hamana, Oguri, Luo, Nishizawa, Takahashi, Nicola, Osato, Kannawadi, Sunayama, Armstrong, Bosch, Komiyama, Lupton, Lust, MacArthur, Miyazaki, Murayama, Nishimichi, Okura, Price, Tait, Tanaka, and Wang, 2023; Dalal, Li, Nicola, Zuntz, Strauss, Sugiyama, Zhang, Rau, Mandelbaum, Takada, More, Miyatake, Kannawadi, Shirasaki, Taniguchi, Takahashi, Osato, Hamana, Oguri, Nishizawa, Malagón, Sunayama, Alonso, Slosar, Luo, Armstrong, Bosch, Hsieh, Komiyama, Lupton, Lust, MacArthur, Miyazaki, Murayama, Nishimichi, Okura, Price, Tait, Tanaka, and Wang, 2023; Asgari, Lin, Joachimi, Giblin, Heymans, Hildebrandt, Kannawadi, Stölzner, Tröster, van den Busch, Wright, Bilicki, Blake, de Jong, Dvornik, Erben, Getman, Hoekstra, Köhlinger, Kuijken, Miller, Radovich, Schneider, Shan, and Valentijn, 2021). Such analyses measure the tiny statistical patterns in aligned or correlated shapes of a vast number of galaxies. To do so requires high-accuracy measurements of their shapes that have been extensively validated (for a review, see Mandelbaum 2018). As galaxy samples become larger with surveys such as Euclid [^1] http://sci.esa.int/euclid (Laureijs et al. 2011), the Vera C. Rubin Observatory Legacy Survey of Space and Time [^2] http://www.lsst.org (LSST: LSST Science Collaboration et al. 2009; Ivezić et al. 2019), and the Nancy Grace Roman Space Telescope [^3] https://roman.gsfc.nasa.gov (Roman: Spergel et al. 2015), our systematic uncertainty requirements are more stringent, demanding improved shape measurement techniques.
@@
Non-zero $m$ and $c$ terms can be introduced due to imperfect (i) PSF modeling and (ii) shear estimation. During PSF modeling, sources such as inaccurate characterization of detector effects, impurity in stellar samples, and mismodeling of atmospheric effects can cause the misestimation of PSF models, and the PSF size and ellipticity errors result in $m$ and $c$ respectively. Even if PSF models are accurate, misevaluation of the PSF for individual galaxies during shear estimation can lead to significant multiplicative bias. We will explore this further in this paper. Furthermore, in the low signal-to-noise (S/N) regime, the pixel noise causes prominent “noise bias” (e.g., Hirata and Seljak 2003; Kacprzak et al. 2012; Melchior and Viola 2012; Refregier et al. 2012), because shape estimation is a non-linear function of pixel intensities. Other well-known significant shear biases include “model bias” due to the mismatch of assumed and the actual galaxy light profile during shape estimation (e.g., Kacprzak et al. 2014; Melchior et al. 2010; Voigt and Bridle 2010) and “selection bias” that happens when the galaxy selections depend on shear (e.g., Hirata and Seljak 2003; Bernstein and Jarvis 2002).
@@
In order to account for these biases, modern lensing analyses calibrate the shears that have been estimated. In this step, one measures the shear response $\partial e_{i}/\partial\gamma_{j}$, where the measured galaxy ellipticity ($e_{i}$) responds to an applied shear ($\gamma_{j}$). By determining the shear response for an ensemble of galaxies, one can calibrate the aforementioned biases over a wide range of galaxy populations. In the Shear TEsting Programme (STEP; Heymans et al. 2006; Massey et al. 2007) and the GRavitational lEnsing Accuracy Testing challenges (GREAT; Bridle et al. 2010; Kitching et al. 2013; Mandelbaum et al. 2015), the lensing community has dedicated efforts to develop new unbiased shear estimation and calibration algorithms and mitigate/self-calibrate these biases for the Stage-III surveys to satisfy their systematics budgets such as the Hyper-Suprime Cam Subaru Strategic Program [^4] http://hsc.mtk.nao.ac.jp/ssp/ (HSC: Aihara et al. 2018), the Kilo-Degree Survey [^5] http://kids.strw.leidenuniv.nl/ (KiDS: de Jong et al. 2013) and the Dark Energy Survey [^6] http://www.darkenergysurvey.org/ (DES: The Dark Energy Survey Collaboration 2005). We briefly introduce a few more recent algorithms that are relevant to this paper. The first widely utilized technique is Metacalibration (Sheldon and Huff, 2017; Huff and Mandelbaum, 2017). This method artificially applies a shear to each object and numerically measures the response. It can empirically correct for “noise”, “model”, and “selection” biases to a few percent (Sheldon and Huff 2017; Gatti, Sheldon, and others 2021). Additionally, Bayesian Fourier Domain (BFD;  Bernstein and Armstrong 2014; Bernstein et al. 2016b) can independently self-calibrate the biases by constructing templates from deep-field galaxies and building the probability distribution of Fourier-space moments for wide-field galaxies. We also point out that AnaCal, which computes the shear response of basis functions and propagates the shear response of basic modes to derive the shear response of the final observable, has shown the potential to calibrate shear calibration biases to a sub-percent level (Li and Mandelbaum, 2023, 2024).
@@
With this progress, a number of challenges remain, which motivate the advancements in this work. In PSF modeling, Meyers and Burchat (2015) identified the differential chromatic refraction (DCR) and the wavelength dependence of seeing as primary drivers of PSF misestimation. These chromatic effects will be especially essential to be accounted for in Stage-IV surveys. Further, they have already been detected in previous DES analyses (Jarvis et al. (2021), Gatti, Sheldon, and others (2021); hereafter GS20). In shear estimation, as the number of single-epoch (SE) images increases for a given region on the sky, naive image coaddition introduces SE image edges into the final coadd image, which produce discontinuities in the coadded PSF. These PSF discontinuities can bias subsequent shape measurements performed on the coadd image (Mandelbaum et al., 2023). This effect motivated the techniques that use galaxies detected in coadd images but with measurements taken from the SE images. Unfortunately, this approach has two main limitations. First, the computational cost of the shear measurement increases linearly with the number of SE images. While the DES has only $\sim$10 SE images overlapping a given point on the sky in a given band, Stage-IV surveys will, in some cases, have $\cal{O}$(100) SE images (LSST Science Collaboration et al., 2009). Second, detecting objects is a weak lensing shear-dependent operation, and ignoring this dependence causes significant biases in shear measurements (e.g., Hirata and Seljak 2003; Kannawadi et al. 2019; Sheldon et al. 2020). Finally, as the depths of surveys increase, the fraction of objects that are blended increases as well. The increased blend fraction coupled with detection biases can cause catastrophic biases in shear measurements for future surveys (Sheldon et al. 2020; hereafter S20).
@@
To address these issues, we have pursued a new approach to shear measurement which proceeds in two stages. First, we employ a cell-based coadding algorithm to reprocess the entire DES survey. A cell-based coadd in this work is defined as a small coadd image, approximately one arcminute on a side, within which none of the input SE images have edges. This kind of coadd has a well-defined PSF model (Mandelbaum et al., 2023; Armstrong et al., 2024) and so is suitable for weak lensing shape measurements. We lose some survey depth, unfortunately, in the process of constructing cell-based coadds since not all input SE images fully overlap a given region on the sky (Armstrong et al., 2024). This trade-off is discussed in more detail below. Second, we run the Metadetection weak lensing shear estimation algorithm over the cell-based coadd images. The application of Metadetection on simulated Rubin data by Sheldon et al. (2023) has demonstrated that Metadetection can produce unbiased shear measurements in the presence of detection and blending at Rubin LSST ten-year depths (LSST Science Collaboration et al., 2009) to a fractional accuracy better than $\sim 0.1\%$. While this level of accuracy is not required for DES Y6 cosmological analyses, we aim for $m$ around 1.3% or smaller. This level, according to the LSST DESC Science Requirements Document (The LSST Dark Energy Science Collaboration et al., 2018), would be sufficient for LSST Year 1 analysis and certainly for DES Year 6 analysis. The previous DES Y3 Metacalibration pipeline (Gatti, Sheldon, and others, 2021) produced significant enough biases due to blending to shift the Y3 cosmological contours by $\approx 1\sigma$ (MacCrann, Becker et al., 2022; Amon et al., 2022) if not corrected.
@@
As for additive bias, previous DES analyses have found non-negligible additive bias in DES data (Gatti, Sheldon, and others, 2021) that is beyond cosmic variance and above what was measured in image simulations (MacCrann, Becker et al., 2022). The origin of this bias is unknown, and Amon et al. (2022) chose to subtract it from the data before measuring the shear two-point correlation functions. This procedure is valid under the assumption that the bias is actually constant and when the two-point function measurements are restricted to angular scales where the two-point function amplitudes are significantly larger than the additive signal squared. For non-constant additive signals that are driven by PSF modeling and deconvolution, one can use the data themselves to measure the additive bias signals. For the DES Y3 shear catalog, these biases were shown to be small and could thus be safely ignored. Alternatively, one can use the additive bias measurements to directly marginalize over them during the cosmological analysis (Li et al., 2023). For DES Y6, we will employ similar procedures to account for any residual additive biases, and will also attempt to reduce them using new PSF models with chromatic dependence (Schutt et al., 2025).
@@
For the image simulations, we use an updated version of the DES Y3 image simulation code (MacCrann, Becker et al., 2022), to be presented by Mau, Becker, and others 2025. Finally, this paper is further accompanied by Schutt et al. 2025, which describes the upgraded PSF modeling with Piff (PSFs in the Full FOV; Jarvis et al. 2021) that includes color-dependent modeling. This work reports several diagnostic tests on the models created for Y6 data, showing better PSF models in all grizgriz bands compared to Y3 (Jarvis et al., 2021). Finally, the DES collaboration has a second shear catalogue in preparation using the BFD technique (Gatti, Wetzell, and others, 2025), and this catalogue will eventually act as a strong cross-check on this work.
@@ 2.1 DES Data
  • The SE images: The grizgriz-band calibrated images from the DES processing pipeline, created by DES Data Management (DESDM) (Morganson et al., 2018).
@@ 2.2 Mock Catalogues
In some validation tests of the Metadetection shear catalogue, we require covariance matrices for statistics extracted from complex operations on the data. These covariance matrices need to include both galaxy shape noise and the underlying sample variance contributions from large-scale structure. For these kinds of tests, mock shape catalogues were generated from the shear maps in gravity-only N-body simulations. We make use of CosmogridV1 (Fluri et al., 2022). The dataset is described in Kacprzak et al. (2023). We use 200 independent realisations at the fixed cosmology ($\Omega_{\rm m}=0.26$, $\sigma_{8}=0.84$, $\Omega_{\rm b}=0.0493$, $n_{\rm s}=0.9649$, $h=0.673$). The simulations provide 69 full-sky lens planes between redshift 0 and redshift 3.5 with HEALPix [^8] http://healpix.sourceforge.net (Górski et al., 2005) nside=2048; the lens planes are converted into convergence planes under the Born approximation (e.g., Eq. 2 from Fosalba et al. 2015), and shear planes are obtained from the convergence maps using a full-sky generalisation of the Kaiser and Squires (1993) algorithm (Chang, Pujol, Mawdsley, Bacon, Elvin-Poole, Melchior, Kovács, Jain, Leistedt, Giannantonio, Alarcon, Baxter, Bechtol, Becker, Benoit-Lévy, Bernstein, Bonnett, Busha, Carnero Rosell, Castander, Cawthon, da Costa, Davis, De Vicente, DeRose, Drlica-Wagner, Fosalba, Gatti, Gaztanaga, Gruen, Gschwend, Hartley, Hoyle, Huff, Jarvis, Jeffrey, Kacprzak, Lin, MacCrann, Maia, Ogando, Prat, Rau, Rollins, Roodman, Rozo, Rykoff, Samuroff, Sánchez, Sevilla-Noarbe, Sheldon, Troxel, Varga, Vielzeuf, Vikram, Wechsler, Zuntz, Abbott, Abdalla, Allam, Annis, Bertin, Brooks, Buckley-Geer, Burke, Carrasco Kind, Carretero, Crocce, Cunha, D’Andrea, Desai, Diehl, Dietrich, Doel, Estrada, Fausti Neto, Fernandez, Flaugher, Frieman, García-Bellido, Gruendl, Gutierrez, Honscheid, James, Jeltema, Johnson, Johnson, Kent, Kirk, Krause, Kuehn, Kuhlmann, Lahav, Li, Lima, March, Martini, Menanteau, Miquel, Mohr, Neilsen, Nichol, Petravick, Plazas, Romer, Sako, Sanchez, Scarpine, Schubnell, Smith, Smith, Soares-Santos, Sobreira, Suchyta, Tarle, Thomas, Tucker, Walker, Wester, Zhang, and DES Collaboration, 2018; Jeffrey, Gatti, Chang, Whiteway, Demirbozan, Kovacs, Pollina, Bacon, Hamaus, Kacprzak, Lahav, Lanusse, Mawdsley, Nadathur, Starck, Vielzeuf, Zeurcher, Alarcon, Amon, Bechtol, Bernstein, Campos, Rosell, Kind, Cawthon, Chen, Choi, Cordero, Davis, DeRose, Doux, Drlica-Wagner, Eckert, Elsner, Elvin-Poole, Everett, Ferté, Giannini, Gruen, Gruendl, Harrison, Hartley, Herner, Huff, Huterer, Kuropatkin, Jarvis, Leget, MacCrann, McCullough, Muir, Myles, Navarro-Alsina, Pandey, Prat, Raveri, Rollins, Ross, Rykoff, Sánchez, Secco, Sevilla-Noarbe, Sheldon, Shin, Troxel, Tutusaus, Varga, Yanny, Yin, Zhang, Zuntz, Abbott, Aguena, Allam, Andrade-Oliveira, Becker, Bertin, Bhargava, Brooks, Burke, Carretero, Castander, Conselice, Costanzi, Crocce, da Costa, Pereira, De Vicente, Desai, Diehl, Dietrich, Doel, Ferrero, Flaugher, Fosalba, García-Bellido, Gaztanaga, Gerdes, Giannantonio, Gschwend, Gutierrez, Hinton, Hollowood, Hoyle, Jain, James, Lima, Maia, March, Marshall, Melchior, Menanteau, Miquel, Mohr, Morgan, Ogando, Palmese, Paz-Chinchón, Plazas, Rodriguez-Monroy, Roodman, Sanchez, Scarpine, Serrano, Smith, Soares-Santos, Suchyta, Tarle, Thomas, To, Weller, and DES Collaboration, 2021). Mock DES Y6 shear fields are obtained by integrating the shear planes assuming preliminary DES Y6 redshift distributions; last, we create mock shear catalogues by sampling the simulated shear fields at the DES Y6 Metadetection galaxy positions and adding DES Y6 shape noise. This last step is performed by randomly rotating the DES Y6 galaxy ellipticities and adding them to the simulated shear field sampled at each galaxy position. Since we can cut four DES Y6 footprints out of each simulation, this procedure provides 800 independent shear catalogues that can be used to infer covariance matrices and perform statistical tests.